OpenStudy (anonymous):

A cyclist rides his bike at a speed of 27 miles per hour. What is this speed in feet per second? How many feet will the cyclist travel in 10 seconds? In your computations, use the fact that 1 mile is equal to 5280 feet. Do not round your answers.

OpenStudy (jdoe0001):

\(\bf \cfrac{27mi}{hr}\implies \textit{27 miles per 1 hour}\\ \quad \\ \cfrac{5280ft}{mi}\quad or\quad \cfrac{mi}{5280ft}\implies \textit{5280 feet in 1 mile}\qquad thus\\ \quad \\ \cfrac{27\cancel{mi}}{\cancel{hr}}\cdot \cfrac{5280ft}{\cancel{mi}}\cdot \cfrac{\cancel{hr}}{60\cancel{mins}}\cdot \cfrac{\cancel{mins}}{60secs}\)

OpenStudy (anonymous):

now what

OpenStudy (jdoe0001):

now that's it :)

OpenStudy (jdoe0001):



OpenStudy (anonymous):

thats not the answer

OpenStudy (anonymous):